OpenStudy (anonymous):

Let f(x)=3+×^2+tan((pi)(x)/2), where -1 <×<1 A) find f^-1 (3) B) find f (f^-1 (5))

OpenStudy (anonymous):

I know that when you inverse a function you switch the x and the y values. But I don't know what to do with the y value instead of x in tangent.

OpenStudy (rational):

for part A, you need to solve : \[3 = 3+x^2+\tan(\pi x/2)\]

OpenStudy (anonymous):

Wouldn't the x value be 3?

OpenStudy (rational):

nope, as the name says, \(f^{-1}(3)\) is the value of \(x\) that makes the function \(f(x)\) spit out \(3\)

OpenStudy (rational):

\[f^{-1}(3) = a \implies f(a) = 3\] so you need to solve \(f(a)=3\) for \(a\)

OpenStudy (anonymous):

Ohhhh

OpenStudy (rational):

\[3 = 3+x^2+\tan(\pi x/2)\] \[0=x^2+\tan(\pi x/2)\] you can eyeball the value of \(x\) that satisfies above equation

OpenStudy (rational):

Yes!

OpenStudy (rational):

so \(f(0) = 3 \implies f^{-1}(3) = 0\)

OpenStudy (anonymous):

Wow so much simpler than I was doing it before! Thank you

OpenStudy (rational):

np what about part B

OpenStudy (rational):

hold up, we don't really need to find \(f^{-1}(5)\)

OpenStudy (anonymous):

And then solve for the inverse and then place into the original

OpenStudy (rational):

\(f\) eats \(f^{-1}\)

OpenStudy (rational):

\[\require{cancel} \large { f(f^{-1}(5))\\~\\ \cancel{f}(\cancel{f^{-1}}(5))\\~\\5 }\]

OpenStudy (anonymous):

Are you sure they cancel?

OpenStudy (ikram002p):

they dont cancel they generate new function f(f^-1(x))=x

OpenStudy (rational):

Think of it like this : \(f^{-1}(5)\) is the value of \(x\) that makes \(f(x)\) spit out \(5\), so if you feed that value to \(f(x)\) again, it simply spits out \(5\)

OpenStudy (rational):